OpenStudy (anonymous):

Secx+cscx=1 Prove the identity Will give medal

OpenStudy (lochana):

let's take left hand side and prove that it is actually 1. so left h side is Secx+cscx and multiply it by sinX.. so we get tanX + 1

OpenStudy (lochana):

Are you sure this question is correct?

OpenStudy (lochana):

your question is wrong unfortunately.

OpenStudy (anonymous):

Yes it's correct

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

what's the reciprocal of sec and csc ?

OpenStudy (anonymous):

Cos tan and sin

Nnesha (nnesha):

hmm sec = 1/cos not just cos and csc = 1/sin

OpenStudy (anonymous):

Srry I wrote it wrong

Nnesha (nnesha):

\[\frac{ 1 }{ \sin x } +\frac{1}{\cos x} =1\] now find common denominator

Nnesha (nnesha):

or you can do the other way like lochana mentioned but when we prove identity we should work only one side (most of the case where we only work on one side )

Nnesha (nnesha):

on*

OpenStudy (anonymous):

So work on the left side

OpenStudy (lochana):

@Nnesha yes. You are correct. But I don't think you can make it. because the only way that secx+ cscx =1 when x = 0

Nnesha (nnesha):

correct \[secx +\csc x \cancel{ =}1\]

Nnesha (nnesha):

multiply by sin(we should multiply *both* sides by sin)\[\rm sinx(\frac{ 1 }{ \sin x } +\frac{1}{\cos x}) =1*sinx\] sorry i meant to say multiply by sin

OpenStudy (anonymous):

I multiply by tan

Nnesha (nnesha):

distribute parentheses by sin

Nnesha (nnesha):

by sin** not tan sorry that was a mistake

OpenStudy (anonymous):

Cscx

Nnesha (nnesha):

what about it ??

OpenStudy (anonymous):

Is that what it equals

Nnesha (nnesha):

hmm no distribute parentheses by sin x \[\rm \color{Red}{sinx}(\frac{ 1 }{ \sin x } +\frac{1}{\cos x}) =1*sinx\] \[\rm \color{red}{sinx}*\frac{1}{sinx}+\color{ReD}{sinx}*\frac{1}{cosx}\] simplify left side

OpenStudy (anonymous):

1+tan

Nnesha (nnesha):

yes right that's how lochana got tanx+1 at left side \[\rm tanx+1=sinx\] both sides aren't equal

OpenStudy (anonymous):

Oh okay so my question wouldn't be an identity because they don't equal

OpenStudy (lochana):

@bluenile955 you are correct. you end up with getting 1+ tanx . which is only give you 1 when x = 0 .

OpenStudy (anonymous):

Thx